OpenStudy (anonymous):

2 friends worked out on treadmills at the gym John walked 2 miles in 2/3 hour! Julian walked 1 1/3 in 30 minutes! Who walked at a faster rate? Explain

rebeccaxhawaii (rebeccaxhawaii):

well 2/3 of an hour is 40 minutes

OpenStudy (mathmale):

Find the unit rate for each. John walked ? miles in 1/2 hour .... so how many miles in 1 full hour? Julian walked ? miles in how many hours? You could also 'forget about hours' and use 'minutes' as your unit of time. Please show all work. Last, we'll compare the unit rates of the two men and identify which is the greater (faster walker).

OpenStudy (texaschic101):

if it was 2 miles in 40 min, then it is 1 mile in 20 min...or 3 miles per hr. 1 1/3 miles in 30 min, in an hr, it is doubled...4/3 * 2 = 8/3 = 2 2/3 miles per hr. does that help

OpenStudy (mathmale):

Who walks faster?